OpenStudy (anonymous):

please help me! find the variables: x+2y-z=4 3x-y+2z=3 -x+3y+z=6

OpenStudy (anonymous):

1. Substitution 2. Gaussian Elmination 3. Cramer's Rule

OpenStudy (anonymous):

First, to solve for y, add the first and third equations. x + 2y - z = 4 -x + 3y + z = 6 5y = 10 y = 2 Now plug it back in. x + 4 - z = 4 x - z = 0 3x - 2 + 2z = 3 3x + 2z = 5 x - z = 0 3x + 2z = 5 Now can you solve this system of linear equations?

OpenStudy (anonymous):

yes thank you for explaining!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

@Logics, you may find it useful to look up the other two methods I mentioned. :) They both have their drawbacks and upsides, just as substitution does.

OpenStudy (anonymous):

ok thank you :)